See Candy had a FCFE of $6.1M in 2016 and has 2.32M shares outstanding. See's cost of equity (required return) is 9.3%. If FCFE is expected to grow at 6.5% forever, the price per share of See Candy should be ____________.

A- $108.00

B- $68.29

C- $100.01

D- $232.02
Answer #1

HI

Here per share FCFE = 6.1/2.32 = $ 2.63

cost of equity k = 9.3%

growth rate g = 6.5%

So Price per share P = per share FCFE*(1+g)/(k-g)

   =2.63 *(1+6.5%)/(9.3%-6.5%)

   = $100.01

Thanks